Consider a production line with three single-machine stations in series. Each has processing times with mean two hours and standard deviation of two hours.

a. Suppose we run this line as a push system and release jobs into it at a rate of 0.45 per hour with arrival variability given by Ca = 1. What is the average WIP in the line?

b. Compute the throughput of this line if it is run as a CONWIP line with a WIP level equal to your answer in (a). What is the throughput of the CONWIP line?
Answer #1

Processing time (Mean) = 2 hours

Standard deviation = 2 hours

Rate of jobs = 0.45 per hours

Utilization Server (u) =

= 0.015 per hour

Time spent in a station will be equal to

Given, Ca = 1

Thus, Time spent in a station (CT) = 2.091 hours

(a) Average WIP = CT * TH

= 2.091 * (2/60) = 0.068 Jobs

(b) TH = 2/60

= 0.033